OpenStudy (ffasinger):

A circle has an area of 36 m2. What is the area of a 40 sector of this circle? A. 6 m2 B. 4 m2 C. 2 m2 D. 3 m2

OpenStudy (ffasinger):

@Directrix

OpenStudy (anonymous):

40 sector?

OpenStudy (anonymous):

Do you have a picture?

OpenStudy (wolf1728):

Should that be 40 degree sector?

OpenStudy (ffasinger):

No image on this one.

OpenStudy (anonymous):

@wolf1728 @FFASinger Assuming it's 40 degrees. Then: 40/360 = 0.11111 0.111111111 * 36 = 4m^2

Directrix (directrix):

Here is an illustration of the sector and finding its area. This problem seems easier than the last one

Directrix (directrix):

40/360 is the circle's area swept out by the sector. 360 degrees in a circle. The sector has 1/9 of the circle's area. (Note: 40/360 = 1/9) If the circle's area is given to be 36, then the sector is occupying 1/9 of that area. Multiply ( 1/9 * 36 ) to get the area of the sector with central angle 40. @FFASinger